1/4 of a certain number is added to 4 1/3,The sum is the same as when 1/3 on it is subtracted from 20 2/3.find the number
Answer by Alan3354(69443)

1/4 of a certain number is added to 4 1/3,The sum is the same as when 1/3 on it is subtracted from 20 2/3.find the number
-------
n = the number
n/4 + 13/3 = 62/3 - n/3
Multiply by 12
3n + 52 = 248 - 4n
7n = 196
n = 28